# Mapping the Spiral of Silence: Surveying Unspoken Opinions in Online Communities

### v1

@@ 4.1 Topic Generation
Ultimately, from the 33 topics, we randomly selected 11 for our final study to ensure there are an adequate number of responses per topic for our analysis. These topics covered a range of political and social issues, including abortion rights, military spending, and affirmative action. Each identified topic contains two viewpoints that articulate two main clusters of opinion on the topic: for example, with affirmative action, one viewpoint states, “I believe affirmative action counters systemic biases and fosters a diverse, inclusive society,” and the other viewpoint states, “I believe affirmative action could unintentionally cause reverse discrimination and undermine merit, potentially increasing societal division.” For the full list of topics and viewpoints, refer to Appendix Table 4.
4.2 Opinion Expression Survey
Next, we use the generated viewpoints to gauge the rate of self-silencing among active community members. Since these self-silenced viewpoints will not be posted on Reddit, we directly survey community members to measure this phenomenon. The goal of our survey is to identify how willing an individual is to share their opinion on a topic to a subreddit. Specifically, we are interested in the following dimensions: (1) differences in sharing behavior when individuals view themselves as being in the majority versus minority; and (2) how self-silencing differs across subreddits.
Survey Instrument
For each topic, our survey captures the likelihood that participants would share their viewpoint on each subreddit. After selecting subreddits and passing the screening quiz for those subreddits, participants chose up to five topics they thought were relevant to each subreddit. For each topic, they indicated the viewpoint they agreed with. Importantly, this process allowed us to validate that the generated topics are germane to the subreddit. For example, for r/combatfootage, selected topics included “ethics of drone warfare in the Middle East,” “military spending,” and “the role of NATO in maintaining global peace.” In r/askfeminists, the only relevant topic was “abortion rights.”
After selecting their preferred viewpoint, participants rated the extent to which they agreed with the viewpoint, how much they believed the majority of subreddit members agreed with the viewpoint, and whether they thought the viewpoint should be allowed to be posted on the subreddit. For all viewpoints the participant agreed with and thought should be allowed on the subreddit, we asked about their likelihood of sharing that viewpoint on the subreddit and their likelihood of upvoting that viewpoint if someone else shares it on the subreddit. The order in which the topics were presented for each subreddit was randomized.
Participants
Survey participants were recruited via Prolific. To qualify, participants were required to live in the United States, be at least 18 years old, and identify as an active member of at least one political subreddit included in our list. This list contains r/politics and r/worldnews as well as the 21 additional political subreddits from Section 4.1. We include these additional subreddits to obtain a larger pool of participants.
Based on an a priori power analysis, we needed to collect 624 responses (participants $\times$ topics $\times$ subreddit), or in other words, approximately 124 participants to achieve $90\%$ power if $\alpha=0.05$ and a change in $R^{2}>0.02$ were assumed to be the minimum effect size of interest. We collected 906 responses from 125 participants. Since we focus on political topics, we recruited a balanced sample of participants by political leaning. We asked for participant’s political leaning in our screening quiz and stratified recruitment, resulting in 40 Democrats, 38 Republicans, 40 Independents, and 7 participants with other political ideologies. For quality assurance, we removed 46 responses from 10 participants who failed our attention checks. We also removed individual responses if: (1) the participant felt the viewpoint should not be shared on the subreddit (N=74); (2) the participant selected “don’t know” when asked to pick a viewpoint for a topic (N=122); (3) the participant later disagreed with the viewpoint they selected as representing their stance (N=31); (4) the participant selected a subreddit that we later did not have adequate data to analyze (N=6). After filtering, we used the remaining 627 responses (i.e., likelihood of sharing a viewpoint for a specific subreddit) covering 33 subreddits for our analysis, meeting our original power analysis requirements.
4.3 Measures
Likelihood of Opinion Expression
Our main dependent variable is a participant’s likelihood of opinion expression. Participants are asked to imagine they see a post on their selected subreddit related to the controversial topic. Then we ask their likelihood of sharing the viewpoint, or Share Likelihood, from their main account on a scale from 1 to 7. We also ask participants to provide the likelihood of upvoting the viewpoint (Upvote Likelihood) on a seven-point scale.
Opinion Incongruency
For a given viewpoint, Incongruency is a binary variable indicating whether a participant believes their opinion is held by the majority (0) or minority (1) of other subreddit members. Following prior work (Neuwirth, Frederick, and Mayo 2007; Chia 2014), we operationalize incongruency as the difference between how much a participant agrees with a viewpoint and how much they believe the majority of subreddit members agree. Both agreement scores are measured on a 7-point scale. When converting the difference into a binary variable, we indicate all instances when the participant agrees with the viewpoint and they believe the majority of subreddit members also agree as 0. Conversely, if the participant agrees with the viewpoint but believes the majority of subreddit members are neutral or disagree, incongruency is marked as 1.
Community Inclusion and Diversity
We collect self-perceived community inclusivity and diversity from participants using Weld, Zhang, and Althoff (2022)’s instrument for measuring community values across subreddits. For a given subreddit, participants report their perception of the current state of Inclusion (i.e., “How included and able to contribute do new and existing members feel?”) and Diversity (i.e., “How diverse are the people?”). The response was recorded on an 11-point scale (from 1-11) for consistency with prior work (Weld, Zhang, and Althoff 2022).
Community Moderation
We use content removal as our main measure of community moderation. Using Pushshift data from Jan. 2022 - Dec. 2023, we compute Content Removal Rate as the percentage of submissions removed by moderators within a subreddit (Jhaver, Bruckman, and Gilbert 2019). In total, there were 1.74 million submissions across 33 subreddits. The removal rate ranged from $4.1\%$ (r/austrianeconomics) to $46.0\%$ (r/againsthatesubreddits), with a median of $26.3\%$ percent per subreddit. We apply a logarithmic transformation with base 2 after adding a start-value of 1 to the computed values.
Controls
In addition to binary variables for demographics (gender, race, and political leaning), we include the following controls for personality and Reddit usage:
  • Willingness to self-censor (WTSC): Individual personality differences can impact the likelihood of sharing (Gearhart and Zhang 2014). We use Hayes, Glynn, and Shanahan (2005) well-validated composite measure, consisting of eight questions capturing an individual’s likelihood of withholding their opinion when they sense others disagree. A larger WTSC value means the individual is more disposed to self-censoring.
  • Agreement Intensity: How strongly an individual agrees with a viewpoint impacts the chance that the spiral of silence occurs (Matthes, Rios Morrison, and Schemer 2010). We measure agreement on a 7-point scale from “strongly disagree” to “agree.” Since we only consider viewpoints that participants agree with, we report intensity on a 3-point scale from “somewhat agree” to “strongly agree.”
  • Posting: To control for different base rates of posting frequency (Gong, Lim, and Zhu 2015), we include a measure, Posting, which captures a participant’s self-reported likelihood of posting (over lurking) on a 5-point scale.
  • Account Tenure: Since newcomers may exhibit different sharing behavior compared to platform veterans (Yang, Kraut, and Levine 2017), we include a control capturing the age of the participant’s Reddit account in years.
5 Results
Figure 2: Across all topics, the average percentage of incongruent viewpoints likely to be shared is $28.3\%$. On the left, we report the percentage of incongruent viewpoints willing to be shared (i.e., Share Likelihood $>$ 4) out of all incongruent viewpoints for the given topic. On the right, we do the same for congruent viewpoints.
In this section, we analyze the responses from our survey using linear mixed-effect models. We start by presenting the model we used for analysis. Next, we compare opinion expression in congruent versus incongruent opinion climates. Then, we explore the impact of community-level factors, such as inclusivity and moderation practices. We conclude by analyzing what factors modulate upvoting behavior.
5.1 Method of Analysis
To analyze the results of our survey, we use linear mixed-effect models, with subreddit as a random effect. [^1] We report our analysis with ordinal regression models in Appendix B and find qualitatively similar results. We introduce a series of three nested models. As shown in Table 1, we first have a model using only our control variables as predictors. We then add additional measures for viewpoint incongruency, community inclusivity, diversity, and moderation. Finally, we include interaction effects between incongruency and our measurements for community values and moderation to our model. We standardize all continuous independent variables and controls. For our first set of models (Models 1a-1c), the dependent variable is the self-reported likelihood of sharing the viewpoint on a main account (Share Likelihood). In Models 2a-2c, we use the likelihood of upvoting the viewpoint (Upvote Likelihood) as our dependent variable. The dependent variables are left in raw units.
Table 1: Although incongruent viewpoints are less likely to be shared than congruent viewpoints, having more inclusive communities can help abate self-silencing behavior. We present the results of a linear mixed-effect model that predicts a participant’s likelihood of sharing their opinion as a hierarchical regression. Model 1a contains only our controls. Model 1b includes the controls and fixed effects for opinion incongruency, community values, and community moderation. Model 1c adds interaction effects. Note: p$<$0.05; ∗∗p$<$0.01; ∗∗∗p$<$0.001.
Model 1a Model 1b Model 1c
Coef. SE $p$ Coef. SE $p$ Coef. SE $p$
(Intercept) $4.27$ $0.25$ $0.00^{***}$ $4.56$ $0.27$ $0.00^{***}$ $4.58$ $0.27$ $0.00^{***}$
Controls
WTSC $-0.38$ $0.08$ $0.00^{***}$ $-0.37$ $0.08$ $0.00^{***}$ $-0.38$ $0.08$ $0.00^{***}$
Agreement Intensity $0.44$ $0.08$ $0.00^{***}$ $0.35$ $0.08$ $0.00^{***}$ $0.37$ $0.08$ $0.00^{***}$
Male (0/1) $-0.41$ $0.18$ $0.03^{*}$ $-0.47$ $0.18$ $0.01^{**}$ $-0.48$ $0.18$ $0.01^{**}$
White (0/1) $-0.05$ $0.17$ $0.79$ $0.03$ $0.17$ $0.85$ $0.04$ $0.17$ $0.79$
Democrat (0/1) $-0.40$ $0.17$ $0.02^{*}$ $-0.65$ $0.17$ $0.00^{***}$ $-0.64$ $0.17$ $0.00^{***}$
Account Tenure $0.25$ $0.08$ $0.00^{***}$ $0.20$ $0.08$ $0.01^{*}$ $0.20$ $0.08$ $0.01^{*}$
Posting $0.51$ $0.08$ $0.00^{***}$ $0.39$ $0.08$ $0.00^{***}$ $0.40$ $0.08$ $0.00^{***}$
Independent Variables
(H1) Incongruency (0/1) $-0.73$ $0.18$ $0.00^{***}$ $-0.63$ $0.19$ $0.00^{***}$
(H2a) Inclusion $0.38$ $0.08$ $0.00^{***}$ $0.28$ $0.10$ $0.00^{**}$
(H2b) Diversity $0.06$ $0.08$ $0.49$ $0.06$ $0.09$ $0.51$
(H3) Content Removal Rate $0.30$ $0.17$ $0.08$ $0.31$ $0.16$ $0.07$
(H2a) Incongruency $\times$ Inclusion $0.35$ $0.18$ $0.05^{*}$
(H2b) Incongruency $\times$ Diversity $-0.01$ $0.18$ $0.97$
(H3) Incongruency $\times$ Content Removal Rate $-0.16$ $0.19$ $0.39$
Marginal $R^{2}$ $0.17$ $0.22$ $0.22$
Conditional $R^{2}$ $0.27$ $0.35$ $0.34$
Figure 3: An increase in community inclusion has a greater positive impact on the likelihood of sharing incongruent viewpoints compared to congruent ones. We visualize the marginal effects of community inclusion on opinion expression for incongruent and congruent viewpoints. For every one standard-deviation unit increase in perceived inclusion, the likelihood of sharing increases by $0.63$ for incongruent viewpoints and $0.28$ for congruent viewpoints.
5.2 Analyzing Opinion Sharing Behavior
Participants are slightly likely to share congruent viewpoints.
We start by examining the relationship between the likelihood of sharing a viewpoint and our control variables. In Model 1a, the intercept is $4.27$, representing the likelihood of sharing a congruent viewpoint for a non-Male, non-White, and non-Democrat individual who posts on Reddit rarely to occasionally, has an account age of 5.8 years, and WTSC of 3.68 out of 8. The intercept value falls between “neither likely nor unlikely” and “slightly likely” on the 7-point scale from the survey.
Incongruent viewpoints are shared less often.
Participants were less likely to voice incongruent viewpoints across all topics despite agreeing with the stance. Approximately half ($51.5\%$) of participants responded that they are likely to share their viewpoint (Share Likelihood $>4$) when they believe themselves to agree with the majority (see Fig. 2). In contrast, the proportion of participants likely to share incongruent viewpoints is lower, ranging from $11\%$ to $42\%$. On average, across topics, participants are twice as likely to share a viewpoint they perceive as being in the majority compared to viewpoints they believe are in the minority.
We also examine Model 1b, which includes a fixed effect for viewpoint incongruency, allowing us to compare the likelihood of sharing a viewpoint that is perceived to be in the majority versus minority. In line with the spiral of silence theory (Noelle-Neumann 1974), participants are less likely to comment their viewpoint when they believe their opinion differs from that of the majority ($p<0.001$). The likelihood of sharing decreases by $-0.73$ for an incongruent viewpoint compared to a congruent viewpoint. On average, when participants perceive themselves as being in the minority, they are less likely to share their viewpoint ($M=3.01$) compared to those who view themselves as in line with the majority ($M=4.05$). While participants view themselves as being congruent with the majority for most viewpoints, $72.2\%$ (N=78) of participants hold at least one incongruent viewpoint. Given that users must self-select into being members of different communities on Reddit, it is unlikely that they will be completely incongruent across all topics. Although community members are not typically in the opinion minority, they are likely to feel incongruent on at least a small selection of topics discussed on the subreddit.
Community inclusivity positively influences sharing behavior.
Model 1b also includes measures for perceived community inclusion and diversity. Inclusion has a positive relationship with Share Likelihood ($p<0.001$) — the likelihood of sharing increases by $0.38$ for each standard-deviation increase in perceived community inclusivity. In other words, when participants perceived a subreddit as being more open to contributions and felt more included in the community, they were more likely to participate.
Furthermore, we find evidence supporting H2a: that the positive relationship between inclusivity and viewpoint sharing is more pronounced for incongruent opinions. Model 1c includes an interaction effect between Incongruency and Inclusion. The main effect of inclusion continues to be significant ($p<0.01$), and there is a modest but positive relationship between the interaction effect and the likelihood of sharing ($\beta=0.35$, $p=0.05$). For an incongruent viewpoint, a one-standard-deviation increase in inclusivity results in the sharing likelihood increasing by $0.63$; for congruent viewpoints, the sharing likelihood only increases by $0.28$ (see Fig. 3). This suggests that fostering more inclusive communities can help combat the spiral of silence.
However, we do not find evidence to support H2b, that perceived community diversity is more positively related to opinion sharing for incongruent viewpoints ($p=0.97$). It is possible that in politics-oriented subreddits, diversity is less of a concern to users. When comparing the relative importance of different values across subreddits, Weld, Zhang, and Althoff (2024) found that communities focused on news, such as those included in our analysis, placed the lowest priority on diversity.
Content removal has no-to-weak association.
We do not find evidence at our level of statistical power to support a claim that more moderation is associated with a greater likelihood of opinion-sharing. The percentage of content removed has a positive but non-significant trend with opinion expression ($\beta=0.30$, $p=0.08$). We also do not find any significant interaction between Incongruence and Content Removal Rate ($p=0.39$).
Finally, we explored whether there is a quadratic relationship between moderation and participation, since draconian moderation policies could stifle people’s desire to speak up (Kraut and Resnick 2012) (see Appendix B for full results). While there may be weak evidence of such trend ($\beta=-1.10$), the relationship is non-significant ($p=0.13$). It is possible that the subreddits in our sample are not overly moderated to the point that participation is diminished or subreddit members are unaware of the extent of content removal within a community (Jhaver et al. 2019).
Table 2: Incongruent viewpoints are less likely to be upvoted compared to congruent viewpoints. We present the results of a linear mixed-effect model that predicts the likelihood of upvoting as a hierarchical regression. Model 2a only includes the controls; Model 2b has the controls and fixed effects for Incongruency, community values, and community moderation; and Model 2c adds interaction effects. Note: p$<$0.05; ∗∗p$<$0.01; ∗∗∗p$<$0.001
Model 2a Model 2b Model 2c
Coef. SE $p$ Coef. SE $p$ Coef. SE $p$
(Intercept) $5.37$ $0.19$ $0.00^{***}$ $5.39$ $0.21$ $0.00^{***}$ $5.43$ $0.21$ $0.00^{***}$
Controls
WTSC $0.01$ $0.06$ $0.93$ $0.01$ $0.06$ $0.85$ $-0.00$ $0.06$ $0.94$
Agreement Intensity $0.50$ $0.06$ $0.00^{***}$ $0.46$ $0.06$ $0.00^{***}$ $0.47$ $0.06$ $0.00^{***}$
Male (0/1) $0.01$ $0.13$ $0.96$ $-0.03$ $0.13$ $0.83$ $-0.05$ $0.13$ $0.67$
White (0/1) $-0.06$ $0.13$ $0.66$ $-0.03$ $0.13$ $0.82$ $-0.03$ $0.13$ $0.81$
Democrat (0/1) $0.29$ $0.13$ $0.03^{*}$ $0.19$ $0.13$ $0.14$ $0.19$ $0.13$ $0.14$
Account Tenure $-0.07$ $0.06$ $0.25$ $-0.10$ $0.06$ $0.09$ $-0.10$ $0.06$ $0.10$
Posting $0.49$ $0.06$ $0.00^{***}$ $0.42$ $0.06$ $0.00^{***}$ $0.42$ $0.06$ $0.00^{***}$
Independent Variables
Incongruency (0/1) $-0.27$ $0.13$ $0.05^{*}$ $-0.20$ $0.14$ $0.15$
Inclusion $0.24$ $0.06$ $0.00^{***}$ $0.20$ $0.07$ $0.01^{***}$
Diversity $0.08$ $0.06$ $0.21$ $0.02$ $0.07$ $0.74$
Content Removal Rate $-0.07$ $0.13$ $0.57$ $-0.04$ $0.13$ $0.74$
Incongruency $\times$ Inclusion $0.09$ $0.13$ $0.48$
Incongruency $\times$ Diversity $0.23$ $0.13$ $0.09$
Incongruency $\times$ Content Removal Rate $-0.20$ $0.14$ $0.15$
Marginal $R^{2}$ $0.18$ $0.21$ $0.22$
Conditional $R^{2}$ $0.32$ $0.35$ $0.36$
5.3 Analyzing Upvoting Behavior
We also explore participants’ likelihood of upvoting as another form of opinion expression. Notably, unlike posting or commenting, upvoting is anonymous.
Incongruent viewpoints are less likely to be upvoted, even when people agree with the stance.
We observe a negative relationship between Incongruency and the likelihood of upvoting ($\beta=-0.27$, $p=0.05$ in Model 2b). Similar to sharing behavior, participants are more likely to upvote content in more inclusive communities: the likelihood of upvoting a comment increases by $0.20$ for each standard deviation increase in perceived community inclusion. However, we do not find any significant interaction effects, indicating the measures for community values and moderation do not impact the likelihood of upvoting a congruent viewpoint compared to an incongruent one.
Upvoting provides an alternative for sharing otherwise self-silenced opinions.
Overall, upvoting provides a valuable avenue for users to express their opinion. We compare the number of viewpoints that users report that they are not willing to comment on but are willing to upvote. In congruent conditions, $68.9\%$ of viewpoints that would not be posted (Share Likelihood $\leq 4$) would be upvoted (Upvote Likelihood $>4$). While this percentage of viewpoints is lower in comparison ($52.6\%$) for incongruent viewpoints, for a majority of viewpoints, participants are still likely to use upvoting as a mechanism for expressing their opinions even when they are unwilling to post them.
6 Discussion
In our work, we seek to measure the extent of self-silencing behavior across different political subreddits. Through our analysis, we find evidence that those who believe their opinion to be in the minority are half as likely to post their viewpoint on Reddit, confirming the spiral of silence theory and reinforcing the distortion that social media data may be applying to public opinion. We also discuss community design factors that can help counteract these silencing effects. A summary of our hypotheses and findings can be found in Table 3.
Confirming and measuring the spiral of silence online.
We find robust evidence that participants are less likely to share their viewpoint when they perceive their opinion to be in the minority. Corroborating Gearhart and Zhang (2018)’s findings, we observe that the rate of self-silencing differs across issue types, although overall participants are consistently less willing to express incongruent viewpoints.
Table 3: Summary of hypotheses and findings. $\checkmark$ indicates the hypothesis is supported. “NS” indicates the hypothesis is not supported.
Hypothesis Result
H1: Likelihood of opinion expression is positively associated with viewpoint congruency.
H2a: The perceived inclusivity of a community is more positively associated with the likelihood of opinion expression when users hold incongruent viewpoints.
H2b: The diversity of community members is more positively associated with the likelihood of opinion expression when users hold incongruent viewpoints. NS
Community inclusivity helps counter the spiral of silence.
We identify community design factors that help counteract silencing effects. Previous work (Neubaum 2022; Wu and Atkin 2018) identified platform-level designs that influence the spiral of silence, but it is difficult for users and moderators to intervene at the platform level compared to the community level. We found that perceived inclusivity has a significant positive impact on community member participation. Moreover, inclusivity plays an outsized role in promoting sharing when individuals consider their opinion to be in the minority. Since moderators have more power to influence the values and norms of online communities, this finding suggests that there are feasible interventions to mitigate self-silencing without having to change platform-level design. For example, although community safety is an important aspect of inclusion (Weld, Zhang, and Althoff 2024), prior work found that most anti-social behavior (e.g., personal attacks, bigotry) remains unmoderated on subreddits (Park, Seering, and Bernstein 2022). Measures such as automated toxicity detectors to help moderators remove harmful content more efficiently can be a fruitful way to reduce self-silencing.
Limitations and future work.
We use self-reported measures of opinion expression to capture the spiral of silence. This approach may lack ecological validity compared to studying sharing behavior directly on the platform. It is possible that participants’ reported willingness to share may differ from their situated behaviors. However, our approach remains well validated in prior work (Gearhart and Zhang 2014, 2018; Chia 2014).One avenue for future work is to build systems that allow us to measure self-silencing directly on the platform rather than using surveys.
In addition, there are limitations with our methodology that can be explored in future studies. We take an automated approach to generating topics and viewpoints. While we reap the benefits of easily scaling the number of topics we can measure, this method does not guarantee coverage of all controversial topics within a community. There may be niche topics to specific online communities that we are not currently measuring. An alternative tactic is to augment the generated topics with crowdsourced viewpoints from community members, although crowdsourced responses may be skewed by the spiral of silence already occurring in the community. Finally, our analyses are centered on politically-oriented communities on Reddit. Results may differ if we looked at other topic areas or other social media platforms, such as Facebook or Twitter. Further exploration on these other communities or platforms can help bolster the generalizability of our findings.
7 Conclusion
In this work, our goal is to measure the extent of the spiral of silence across different online communities. We also seek to identify community design factors that may amplify or mitigate self-silencing effects. Since directly measuring what is left unsaid using social media data is not possible, we propose a new method using LLMs to propose community-specific topics and viewpoints that are likely to be controversial. Then, we survey users’ likelihood of opinion expression, opinion incongruency, and perceived community values. In total, we collected 627 responses from 108 Reddit users across 33 subreddits and 11 topics on their likelihood of opinion expression. Although self-silencing behavior is prevalent, community-level design decisions, such as promoting inclusivity, can encourage opinion sharing.
Appendix A Additional Method Details
A.1 Generating Controversial Viewpoints
Prompts
1You are a member of r/\${SUBREDDIT}. Your task is to generate controversial issues,
2aka issues that will lead to disagreement between r/\${SUBREDDIT} members.
3<name>
4\${SUBREDDIT}
5</name>
6<description>
7\${DESCRIPTION}
8</description>
9<rules>
10\${RULES}
11</rules>
12Provide \${NUMBER} issues that will be controversial in r/\${SUBREDDIT}.
13Return just the issue, no justification.
1You are a member of r/\${SUBREDDIT}. Your task is to generate viewpoints that
2r/\${SUBREDDIT} members would hold on a given topic.
3<name>
4\${SUBREDDIT}
5</name>
6<description>
7\${DESCRIPTION}
8</description>
9<rules>
10\${RULES}
11</rules>
12For each side of the issue, write a 50-word Reddit comment from the perspective of that side that
13follows the rules of r/\${SUBREDDIT}.
14Issue: \${TOPIC}
15Return only the comments as a list, no justification. Example output: ["I am pro-choice.", "I am pro-life."]
Finally, we shorten each of the viewpoints that was generated using the following prompt. We use a temperature of 00 for this task.
1Shorten a statement while retaining the same viewpoint.
2The statement should be phrased as an opinion.
3Text: \${VIEWPOINT}
4Return just the shortened statement, no justification.
Generated Topics and Viewpoints
Table 4: The eleven topics, and their respective viewpoints, used in the opinion expression survey

### v2

@@ 4.1.2. Selecting Political Subreddits
Since we seek to compare variation in opinion expression on topics based on community-level differences, we want to identify a single, static set of topics that are likely to be relevant across all 23 subreddits. To ensure broad applicability of the topics, we selected r/politics and r/worldnews as input to the LLM for topic generation based on their size and generality, rather than using more specialized subreddits that might yield niche topics with limited relevance across communities. Using our pipeline, we generate 40 topics (i.e., 20 per subreddit) with opposing viewpoints for each topic. As shown in Table 1, topics covered a range of political and social issues, including abortion rights, military spending, and affirmative action. Each identified topic contains two viewpoints that articulate two main clusters of opinion on the topic: for example, with affirmative action, one viewpoint states, “I believe affirmative action counters systemic biases and fosters a diverse, inclusive society,” and the other viewpoint states, “I believe affirmative action could unintentionally cause reverse discrimination and undermine merit, potentially increasing societal division.” For the full list of topics and viewpoints, refer to Appendix Table 5.
4.2. Validating Generated Topics and Viewpoints
Once we have our generated topics, how do we ensure that they fall within the bounds of what is acceptable for a given subreddit? We conduct three checks to validate and filter our generations. First, we manually inspect each of the generations for coherence. Then, we survey active community members to certify that the generated topics are relevant and appropriate to the subreddit. Finally, we conduct an “intrusion test” also with active community members to verify that our generated topics could plausibly appear in posts on the subreddit. We apply our checks in a sequential order, meaning that only the topics that pass the coherence check are included in the community-specific relevance check and so on.
4.2.1. Coherence Check
We start by manually filtering the generations for comprehension. Using the 40 topics from the prior step, we check that the generated outputs contained opposing viewpoints, covered controversial topics, and were worded in a manner that participants would understand. From this step, we filtered out seven topics: four had incomplete viewpoints (i.e., the model only provided viewpoints representing one side) and three were either worded confusingly or presented viewpoints that did not present opposite stances, leaving us with 33 topics.
4.2.2. Community-Specific Relevance Check
Next, after ensuring that generations are coherent, we need to validate that our generated topics and viewpoints are not only relevant, but also follow the community norms for the respective subreddit. The first way we do this is by having active members of the subreddits (r/politics and r/worldnews) manually review the topics. We recruited Prolific participants who self-reported as active members. All participants first completed a screening quiz to confirm their knowledge about the subreddits in which they were presented a set of three post titles—two from their selected subreddit and one from another political subreddit (which we manually verify is irrelevant to the selected subreddit)—and asked to select which titles were likely to appear in their selected subreddit (see Appendix A.4 for more details and examples). If they identified all the correct post titles, participants qualified for the survey. In total, 46 participants passed our screening quiz and completed our community-specific topic relevance check. They were then shown a list of generated topics and viewpoints and asked whether they were relevant to the subreddit. Participants also had the option to response “I don’t know” for topics. We removed two topics that many respondents reported not having knowledge on and one topic that was commonly marked as not relevant to the community. Ultimately, from this filtered set of 30 topics, we randomly selected 11 for our final study — five generated with r/politics as the input and six with r/worldnews — to ensure there are an adequate number of responses per topic for our analysis.
4.2.3. Intrusion Testing
While our coherence and community-relevance checks ensure generated topics are understandable and broadly aligned with subreddit norms, they do not tell us whether these topics feel authentic to active community members. A topic may be technically relevant but still strike community members as out of place. To examine whether our topics meet this bar, we conduct an intrusion test, a method of outlier detection widely used in natural language processing as a quantitative evaluation for topic modeling (Chang et al., 2009; Bhatia et al., 2018; Churchill and Singh, 2022).
Task Setup
We recruited a new set of active community members (i.e., Prolific participants who passed our screening quiz for the respective subreddit) to participate in our intrusion testing task. These participants were shown the generated topic and four controversial topics mined from the subreddit that are semantically similar to the generated topic. To obtain the controversial topics mined from the subreddit, we start with Pushshift dumps of all posts made in r/politics and r/worldnews between January 2020 and December 2023 (Baumgartner et al., 2020). Then, following the pipeline from Hessel and Lee (2019), we first discard posts that have less than 30 comments and sort posts by their upvote ratio. After removing posts with a negative score, we randomly select 500 posts in the bottom quartile of scores, which contain topics that subreddit members want to engage with (as indicated by the number of comments under the post) but lead to high amounts of disagreement (as indicated by the close number of upvotes and downvotes.). Finally, we extract common controversial topics using TopicGPT (Pham et al., 2024). For each generated topic, we select the four most semantically similar mined topics as determined by the cosine similarity between sentence embeddings (Reimers and Gurevych, 2019). For example, for the generated topic of “abortion rights”, the mined topics were “impact of Roe v. Wade’s overturning”, “Amy Coney Barett’s confirmation”, “freedom of protest rights”, and “government regulation of personal freedoms and public health.” Participants then identified the “intruder” topic, by selecting which of the five topics is least likely to appear in the subreddit. If our generated topic is inappropriate to a subreddit, then it should be consistently selected as the intruder.
Results
In total, 63 participants completed our task – 32 from r/politics and 31 from r/worldnews. Across the eleven topics, our generated topic was selected as the intruder only 18.8% of the time. A one-proportion $z$-test comparing this intrusion detection rate against random guessing (20.0% selection rate) was not statistically significant (z=0.41z=-0.41, $p=0.68$), indicating that community members were no better than chance at identifying our generated topic. This result suggests that the generated topics are likely to appear in the subreddit according to community members, and for nine of the eleven topics, they are rated as more likely to appear in the subreddit than topics from mined conversation data.
Table 1. Examples of topics, and their respective viewpoints, generated for r/politics and r/worldnews.